OpenStudy (anonymous):

I have a couple questions. Medals will be awarded. Ask me for the questions so I can write the equations.

OpenStudy (anonymous):

The first equation is: \[\left(\begin{matrix}12P ^{3} \\ 15P\end{matrix}\right) \ ^{4}\] Solve using division properties of exponents

OpenStudy (anonymous):

That looks like a matrix to me more than anything else? What exactly is that supposed to be?

OpenStudy (whpalmer4):

Is that supposed to be \[(\frac{12P^3}{15P})^4\]?

OpenStudy (anonymous):

You have to solve using division properties of exponents. You know like \[\frac{ 5^{6} }{ 5^{2} } = 5^{4}\]

OpenStudy (anonymous):

Yes.

OpenStudy (whpalmer4):

Okay, that makes more sense. You can approach this in at least two different ways: 1) simplify the base fraction first, then apply the exponent 2) apply the exponent first, then simplify Which would you like to try?

OpenStudy (anonymous):

Which ever is easier. Personally I would like to do the 1. I have a couple of questions like theese and a few word problems.

OpenStudy (whpalmer4):

Okay, so what is the base fraction after you simplify it?

OpenStudy (anonymous):

I don't know. That is the problem. Is it 0.8 or something bigger?

OpenStudy (whpalmer4):

Well, let's simplify 12/15. What is that as a reduced fraction?

OpenStudy (anonymous):

4/5?

OpenStudy (anonymous):

Right?

OpenStudy (whpalmer4):

Yes. How about P^3/P?

OpenStudy (anonymous):

P^2? Cause you have to subtract. So is it 0.8P^2?

OpenStudy (whpalmer4):

I would stick to the integers. Okay, so you have reduced the base fraction to \[\frac{4P^2}{5}\]If you raise that all to the 4th power, what do you get?

OpenStudy (anonymous):

256P^8/625?

OpenStudy (whpalmer4):

That's correct!

OpenStudy (anonymous):

Awesome. Can I ask you some more questions?

OpenStudy (whpalmer4):

Sure.

OpenStudy (anonymous):

Alright. Next equation.

OpenStudy (anonymous):

\[\left(\begin{matrix}3x ^{2} y ^{5} z ^{-2} \\ 5xz ^{5}\end{matrix}\right) ^{-3}\]

OpenStudy (whpalmer4):

Okay, that's not an equation, technically, as there's no equals sign. It's an expression. First thing I would do is take advantage of the fact that \[a^{-n} =\frac{1}{a^n}\]and for a fraction, we just invert the fraction. That gives us \[(\frac{5xz^5}{3x^2y^5z^{-2}})^3\]Next, I would simplify the base fraction...

OpenStudy (anonymous):

What are the base fractions?

OpenStudy (whpalmer4):

the base fraction is the thing inside the parentheses...that which is raised to the exponent power...

OpenStudy (anonymous):

So raise everything in the parentheses?

OpenStudy (whpalmer4):

But we are going to simplify first...

OpenStudy (anonymous):

I don't know where to start simplifing. I don't know where to start,

OpenStudy (whpalmer4):

You did it for the other one...pick the first letter, simplify that part of the fraction, move on to the next. What are the respective powers of x in numerator and denominator?

OpenStudy (whpalmer4):

If you have the same letter in both numerator and denominator, you can simplify...

OpenStudy (anonymous):

but the x is attached to variables. there is a xz^5 how do you simplify the x's

OpenStudy (whpalmer4):

If you want, you can think of the fraction as \[(\frac{5}{3}*\frac{x}{x^2}*\frac{1}{y^5}*\frac{z^5}{z^{-2}} )^3\]

OpenStudy (anonymous):

(5/3 1/x^1 1/y^5 z^3)^3 (simplified?)

OpenStudy (whpalmer4):

Close. Check your work on the z exponent.

OpenStudy (anonymous):

oh z^7?

OpenStudy (whpalmer4):

Yes. You could write the stuff inside the parentheses (after simplification) as \[\frac{5}{3}x^{-1}y^{-5}z^7\]which makes it easy to now apply the ^3...